# Bibliography.bib

### v1

@@ 1. Introduction
Markov decision processes (MDPs) and reinforcement learning (RL) are widely used mathematical frameworks for decision-making in dynamic and uncertain environments, with an extensive history of research [Bellman, Watkins, RL-Book-Bertsekas.2, RMDP, F.Natural]. In recent years, thanks to the surge in available data and computing power, RL techniques have enjoyed tremendous success across a wide range of applications [Nature1, Nature2, S.levine, Nature3, Autonomous.Driving, Nature4].
Generally, there are two main approaches to RL: model-based [[, see, e.g.,]] Minimax.bound,Kakade.model.based,Li.model_base,OR3 and model-free [[, see, e.g.,]]Bubek,Speedy,SA.Wainwright,VRQ.Wainwright,MDVI,Q-learning-PL. In model-based RL, we first learn a model of the MDP using a batch of state-transition samples and then use this model to form a control policy. On the other hand, model-free approaches directly update either the value function, representing the expected reward starting from each state, or the policy, which is the mapping from states to their actions. Given that model-free RL algorithms operate online, demand less storage space, and are more expressive, the majority of state-of-the-art RL developments have been within the model-free paradigm [Mnih, Mnih2, Trust.Region, Nature3].
@@ 3.1. Cascade Q-Learning: A New Scheme to Reduce the Effect of Noise
Consider an MDP with discount factor $\gamma$ and optimal Q-function $\Theta^{\star}$. Suppose we run Algorithm 1 from the initialization $\Theta_{0}$ for $N_{e}$ iterations with the constant step size $\lambda\in(0,1)$. Then, we have
$\mathbb{E}\|\Theta_{N_{e}}-\Theta^{\star}\|_{\infty}\leq\frac{2\|\Theta_{0}-\Theta^{\star}\|_{\infty}}{(1-\gamma)\lambda N_{e}}+\frac{2}{3}\frac{\gamma}{1-\gamma}\lambda^{2}\log(2D)\|\Theta^{\star}\|_{\text{span}}+\frac{2\lambda}{1-\gamma}\sqrt{2\log(2D)}\big{(}\|\sigma(\Theta^{\star})\|_{\infty}+\sigma_{r}\big{)}.$ (1)
Some comments on Proposition 1 are in order. The first term on the right-hand side of the above inequality indicates the initialization error, while the second and third terms originate from the fluctuations of the Bellman noise $W_{n}=\widehat{\mathcal{T}}_{n}(\Theta^{\star})-\mathcal{T}(\Theta^{\star})$ in the algorithm. A careful look at (1) reveals that the best possible convergence rate for CQ is achieved by choosing a step size of the form $\lambda=\frac{\bar{\lambda}}{\sqrt{N_{e}}}$, where $\bar{\lambda}\in\mathbb{R}_{>0}$ It is interesting to note that $\lambda=\frac{\bar{\lambda}}{\sqrt{N_{e}}}$ is reminiscent of the optimal choice of the step size in the PR averaged stochastic gradient descent algorithm in the optimization literature [[, see, e.g.,]Sec. 2.2]Nemirovski.SA. Hence, Cascade Q-learning seems to be a natural algorithm for addressing fixed problems when the operators exhibit contractive behavior in the $\ell_{\infty}$-norm. Substituting $\lambda=\frac{\bar{\lambda}}{\sqrt{N_{e}}}$ into (1) yields
$\mathbb{E}\|\Theta_{N_{e}}-\Theta^{\star}\|_{\infty}\leq\frac{2\|\Theta_{0}-\Theta^{\star}\|_{\infty}}{\bar{\lambda}(1-\gamma)\sqrt{N_{e}}}+\frac{2\bar{\lambda}^{2}\gamma\log(2D)\|\Theta^{\star}\|_{\text{span}}}{3(1-\gamma)N_{e}}+\frac{2\bar{\lambda}\sqrt{2\log(2D)}\big{(}\|\sigma(\Theta^{\star})\|_{\infty}+\sigma_{r}\big{)}}{(1-\gamma)\sqrt{N_{e}}}.$
It follows directly from the above inequality that by running CQ for
@@ 6.1. Preliminaries
Consider the stochastic processes (16a) and (16b). We denote the fixed points of the Bellman operator $\mathcal{T}$ and the operator $H(\Theta):=\mathcal{T}(\Theta)-\mathcal{T}(\Theta_{m})+\widetilde{\mathcal{T}}(\Theta_{m})$ by $\Theta^{\star}$ and $\widehat{\Theta}$, respectively.
@@ A.2. Proof of Lemma 2
Moreover, using the triangle inequality we find that

### v2

@@ 1. Introduction
Markov decision processes (MDPs) and reinforcement learning (RL) are widely used mathematical frameworks for decision-making in dynamic and uncertain environments, with an extensive history of research [Bellman, Watkins, RL-Book-Bertsekas.2, RMDP, F.Natural, Tac.1997]. In recent years, thanks to the surge in available data and computing power, RL techniques have enjoyed tremendous success across a wide range of applications [Nature1, Nature2, S.levine, Nature3, Autonomous.Driving, Nature4, tac.12].
Generally, there are two main approaches to RL: model-based [[, see, e.g.,]] Minimax.bound,Kakade.model.based,Li.model_base,OR3 and model-free [[, see, e.g.,]]Bubek,Speedy,SA.Wainwright,VRQ.Wainwright,MDVI,Q-learning-PL,TAC.free,modelfree.2. In model-based RL, we first learn a model of the MDP using a batch of state-transition samples and then use this model to form a control policy. On the other hand, model-free approaches directly update either the value function, representing the expected reward starting from each state, or the policy, which is the mapping from states to their actions. Given that model-free RL algorithms operate online, demand less storage space, and are more expressive, the majority of state-of-the-art RL developments have been within the model-free paradigm [Mnih, Mnih2, Trust.Region, Nature3].
@@ 3.1. Cascade Q-Learning: A New Scheme to Reduce the Effect of Noise
Consider an MDP with discount factor $\gamma$ and optimal Q-function $\Theta^{\star}$. Suppose we run Algorithm 1 from the initialization $\Theta_{0}$ for $N_{e}$ iterations with the constant step size $\lambda=\frac{1}{\sqrt{N_{e}}}$. Then, we have
$\mathbb{E}\|\Theta_{N_{e}}-\Theta^{\star}\|_{\infty}\leq\frac{2\|\Theta_{0}-\Theta^{\star}\|_{\infty}}{(1-\gamma)\sqrt{N_{e}}}+\frac{2\gamma\log(2D)\|\Theta^{\star}\|_{\text{span}}}{3(1-\gamma)N_{e}}+\frac{2\sqrt{2\log(2D)}\big{(}\|\sigma(\Theta^{\star})\|_{\infty}+\sigma_{r}\big{)}}{(1-\gamma)\sqrt{N_{e}}}.$ (1)
Some comments on Proposition 1 are in order. The first term on the right-hand side of the above inequality indicates the initialization error, while the second and third terms originate from the fluctuations of the Bellman noise $W_{n}=\widehat{\mathcal{T}}_{n}(\Theta^{\star})-\mathcal{T}(\Theta^{\star})$ in the algorithm. Furthermore, it follows directly from the above inequality that by running CQ for
@@ 6.1. Preliminaries
Consider the stochastic processes (16a) and (16b) with stepsize $\lambda=\frac{1}{\sqrt{N_{e}}}$. We denote the fixed points of the Bellman operator $\mathcal{T}$ and the operator $H(\Theta):=\mathcal{T}(\Theta)-\mathcal{T}(\Theta_{m})+\widetilde{\mathcal{T}}(\Theta_{m})$ by $\Theta^{\star}$ and $\widehat{\Theta}$, respectively.
@@ A.2. Proof of Lemma 2
Moreover, using the triangle inequality, we find that